THE ν_{1} BANDS OF Br^{14}NO_{2} AND Br^{15}NO_{2} AROUND 7.8 μm STUDIED BY HIGH-RESOLUTION FOURIER-TRANSFORM ABSORPTION SPECTROSCOPY

\underline{J. ORPHAL}, Institute of Environmental Physics and Institute of Remote Sensing, University of Bremen, P. O. Box 330440, D-28334 Bremen, Germany; D. SCHEFFLER, B. REDLICH, and H. WILL-NER, Institute of Physical Chemistry and Institute of Anorganic Chemistry, University of Hannover, Callinstr. 9, D-30167 Hannover, Germany.

Nitryl bromide (BrNO_{2}) is a molecule of atmospheric interest. It is probably formed by heterogeneous reactions on polar stratospheric clouds and on sea-salt aerosol particles in the marine troposphere^{a}, and by gas-phase reactions of bromine atoms with NO_{2}^{b}. BrNO_{2} is chemically unstable and difficult to synthesize. From the spectroscopic point of view, BrNO_{2} is a heavy asymmetric top with small rotational constants and some low-energy fundamental vibrations, leading to a rather congested spectrum in the mid-infrared. Recently, we have presented the first high-resolution infrared spectrum of gaseous BrNO_{2} in the region of the b-type ν_{4} bands around 6 μm.^{c} The ground-state rotational constants of BrNO_{2} were determined from the infrared spectra using ground-state combination differences. Although it is possible to estimate the molecular structure using the ground-state rotational constants of ^{79}BrNO_{2} and ^{81}BrNO_{2}, the small displacement of the center-of-mass between these isotopomers causes a strong correlation between the Br-N bond length and the O-N-O angle. Therefore, high-resolution absorption spectra of the a-type ν_{1} bands of Br^{14}NO_{2} and Br^{15}NO_{2} were recorded with a spectral resolution of 0.002 cm^{−1} using the Bruker IFS-120HR FTS at the University of Hannover. The bands were analyzed using a Watson-type A-reduced Hamiltonian in the I^{r} representation^{d}. The analysis confirms the rotational constants derived from line assignments in the ν_{4} bands, and allows for an accurate determination of the ground-state substitution structure of BrNO_{2}.

aB. J. Finlayson-Pitts, F. E. Livingston, and H. N. Berko, Nature 343, 622 (1990)

bW. B. DeMore et al., JPL Publ. 97-4 (1997)

cJ. Orphal, A. Frenzel, H. Grothe, B. Redlich, D. Scheffler, H. Willner, and C. Zetzsch, submitted to J. Mol. Spectrosc. (1998)

dJ. K. G. Watson, in Vibrational Spectra and Structure (Ed. J. R. Durig), 1–89, Elsevier, Amsterdam (1977)